# captionUnknown document class

### v2

@@ VI Conclusion
In this paper, we extended the nominal fall prediction algorithm from [1] to a real-time setting, implementing it in both hardware and simulation. The online implementation yields results similar to the offline version when tested on comparable data. Additionally, in simulation, it achieves an average recovery rate of 0.97. The average runtime in simulation and hardware is approximately 0.0045s and 0.0061s, respectively. The difference of 0.0061s is deemed negligible given the algorithm’s sampling rate of 0.025s (40Hz). Given the similar results between offline simulation and hardware implementations shown in [1], along with the consistent performance between online and offline simulations and between online and offline hardware implementations, and the comparable runtimes between online hardware and simulation implementations, we conclude that offline simulation results are reliable indicators of online performance, and simulation results can effectively predict hardware outcomes.
To further assess the algorithm’s robustness, we evaluated the nominal fall prediction algorithm against a broader range of faults. This evaluation is performed for each of the three components: the critical fault classifier, the lead time classifier, and the lead time regressor. When assessed on this broader range of faults, all three components performed significantly worse—especially for faults in the frontal plane—compared to the results in [1].
To improve the nominal fall prediction algorithm’s performance on this broader fault set while minimizing the need for additional data, we fine-tuned the three components with additional data. The fine-tuned components significantly outperformed the nominal components across all evaluation metrics. Furthermore, the fine-tuned algorithm achieved an average lead time greater than 0.2s with nearly 0 average false positive rate, and a maximum average mean difference of 0.03s for the lead time. These results highlight the fine-tuned fall prediction algorithm’s suitability for real-time and offline fall and lead time prediction in bipedal robots, even across a wide variety of faults.

### v3

@@ VI Conclusion
In this paper, we extended the nominal fall prediction algorithm from [ref:mungai2024fall] to a real-time setting, implementing it in both hardware and simulation. The online implementation yields results similar to the offline version when tested on comparable data. Additionally, in simulation, it achieves an average recovery rate of 0.97. The average runtime in simulation and hardware is approximately 0.0045s and 0.0061s, respectively. The difference of 0.0061s is deemed negligible given the algorithm’s sampling rate of 0.025s (40Hz). Given the consistent performance across online and offline simulations, as well as between online and offline hardware tests, we conclude that offline results are reliable indicators of online performance. Furthermore, the similar results observed between offline simulation and hardware implementations [ref:mungai2024fall], along with the comparable runtimes between online hardware and simulation implementations, suggest that the nominal fall prediction algorithm’s simulation results can effectively predict its hardware outcomes.
To evaluate the nominal fall prediction algorithm’s adaptability to new scenarios, we assessed its robustness against a broader range of faults. This evaluation was conducted for each of its three components: the critical fault classifier, the lead time classifier, and the lead time regressor. All three components exhibited significantly degraded performance under these conditions, particularly for faults in the frontal plane, highlighting key limitations of the nominal algorithm.
To address these limitations while minimizing the need for additional data, we fine-tuned the three components with minimal additional data. The fine-tuned components significantly outperformed the nominal components across all evaluation metrics. Furthermore, the fine-tuned algorithm achieved an average lead time greater than 0.2s, with nearly zero average false positive rate, and a maximum average mean difference of 0.03s for the lead time. These results highlight the fine-tuned fall prediction algorithm’s suitability for real-time and offline fall and lead time prediction in bipedal robots, even across a wide variety of faults.